What is the de Broglie wavelength associated with an electron, accelerated through a potential differnece of 100 volts?

Grade:11

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
KE= 100 eV
So,P=\frac{h}{\lambda }
\lambda =\frac{h}{P}
\lambda =\frac{h}{\sqrt{2mKE}}
Putting the values,
\lambda = 1.227 A^{0}
Thanks and Regards
Apoorva Arora
IIT Roorkee
askIITians Faculty